# Are LLMs Prescient? A Continuous Evaluation using Daily News as the Oracle

### v1

@@ 5 Conclusion and Future Work
We introduce Daily Oracle, a continuously updated QA benchmark leveraging daily news to evaluate the temporal generalization and future prediction capabilities of LLMs. Our experiments reveal that while LLMs maintain a degree of predictive power over future events, their prediction accuracy exhibits a gradual decline over time across various models. Notably, while the stronger model Claude-3.5-Sonnet outperforms others significantly, it still exhibits around 12% performance drop in its post-knowledge cutoff period. Although RAG mitigates the effect of outdated knowledge, a noticeable decline in performance remains. Our findings underscore the necessity for ongoing model updates with more current information and emphasize the importance of disentangling missing knowledge from the lack of up-to-date representations.
We hope this work will draw attention to the need for more practical applications of the continuous training of LLMs, driving advancements in adapting models to real-time data changes. In the future, alongside maintaining Daily Oracle, we plan to incorporate a broader range of models and explore how continuous pre-training and efficient adaptation can address the performance degradation challenges presented in our work.
@@ Acknowledgment
We thank the Microsoft Accelerating Foundation Models Research program for providing Azure cloud compute credits for the GPT APIs. The compute was also supported by the NYU High Performance Computing resources, services, and staff expertise.
@@ Appendix A Dataset Details
A.1 Details for Article Selection
@@
We select daily articles that generate the QA pairs in two ways: (1) Random Selection: We randomly sample three articles each day. (2) Hot Topic Selection: To better capture daily events and reduce noise, we select three articles from the top three hot topics of the day. We identify these hot topics by applying the density-based clustering algorithm DBSCAN (Ester et al.,, 1996) to the new articles based on TF-IDF (Term Frequency-Inverse Document Frequency) representations, forming clusters of news articles for each day. We filter out chaotic clusters by removing those with low average in-cluster cosine similarity scores, which typically correspond to clusters containing a large number of diverse articles. The top three clusters, determined by size, are assumed to represent the most discussed events, i.e. hot topics, since larger clusters indicate more articles covering the same event. One article is picked randomly from each of the top three clusters.
A.2 QA Filtering Principles
Seven principles of QA Filtering step in the data construction process: (1) Correctness of Answers: The answer must be factually accurate and fully aligned with the information in the given article. (2) Non-answerability Before the Publication Date: Since we treat the article’s publication date as the question’s resolution date, the question should not be definitively answerable based on information available before the article’s publication. (3) Absence of Information Leakage: Questions must avoid revealing information that became known only after the article’s publication, maintaining fairness for pre-publication evaluation. (4) Objectivity: Both questions and answers must rely on objective facts, avoiding subjective ideas from the authors. (5) Inclusion of a Clear Temporal Element: Questions must contain a specific and clear reference to time, avoiding vague phrases like “in the future” or “soon.” (6) Public Interest: The questions should address topics of broad public concern. (7) Non-obviousness of the Answer: The answer should not be immediately predictable from the question and must provide new or non-trivial insights.
@@ B.1 Baseline Models Information
B.2 Rejection Rates in the Closed-Book Setting
Figure 8 shows the rejection rates for the closed-book setting. We can see that the rejection rate increases throughout the time for Mistral-7B in TF questions and Mixtral-8x7B in both TF and MC questions. This high rejection rate in TF questions for these two models results in closed-book performance dropping below the random baseline of 50% in certain months, as shown in Figure 4.
Figure 8: Rejection rates for the closed-book setting. We plot the 5-month moving average rejection rates for True/False (TF) and Multiple Choice (MC) questions across different LLMs.
@@ B.2 Rejection Rates in the Closed-Book Setting
To more effectively illustrate the trends of other models at a suitable scale, we display GPT-3.5’s performance in the gold article setting separately. As shown in Figure 9, this outdated model performs relatively poorly throughout. While its accuracy could improve with chain-of-thought prompting (Wei et al.,, 2022), we report its performance using the same prompt format as the other models for consistency in comparison. Nevertheless, a clear downward trend is observed in MC questions.
@@ B.4 More Results in the Constraint Open-Book Setting
Figures 10, 11, 12, 13, 14, 15, and 16 show the constrained open-book evaluation results for more models. Similar patterns are observed as discussed in Section 4.2. Specifically, for Claude-3.5-Sonnet, the constrained open-book performance lags behind its closed-book performance, likely because it already has robust representations of world events, making RAG less effective. Additionally, GPT-3.5 is not included in the constrained open-book setting due to its unexpectedly poor performance in the gold article setting (Figure 9) and budget limitations.
@@
Appendix C Prompts
@@
All the prompts we use are shown in this section. The QA generation prompts and evaluation prompts are adapted from Zhang et al., (2024), and the prompt to categorize our generated questions is taken from Halawi et al., (2024).
@@ 2. Please provide one keypoint that summarizes the new event from the article with the following rules:
Figure 17: Prompt in the Article Summary step, adapted from Zhang et al., (2024).
@@
Figure 18: Prompt in the QA Generation step (part 1), adapted from Zhang et al., (2024).
@@
Figure 19: Prompt in the QA Generation step (part 2), adapted from Zhang et al., (2024).
@@
Figure 20: Prompt in the Misleading Choices Generation step, adapted from Zhang et al., (2024).
@@
Answer: New York officials will implement increased safety precautions for the event. Reasoning: The answer is straightforward and expected, as it directly restates the premise of the question without providing any new or specific details on how the safety precautions will be implemented or what they might entail.
@@
Figure 25: Closed-book evaluation prompt for TF questions, adapted from Zhang et al., (2024).
@@
Figure 26: Closed-book evaluation prompt for MC questions, adapted from Zhang et al., (2024).
@@
Figure 27: Constrained open-book evaluation prompt for TF questions, adapted from Zhang et al., (2024).
@@
Figure 28: Constrained open-book evaluation prompt for MC questions, adapted from Zhang et al., (2024).
@@
Figure 29: Gold article evaluation prompt for TF questions, adapted from Zhang et al., (2024).
@@
Figure 30: Gold article evaluation prompt for MC questions, adapted from Zhang et al., (2024).
@@
Figure 31: Prompt to categorize the generated questions, taken from Halawi et al., (2024).

### v2

@@ 5 Conclusion and Future Work
We introduce Daily Oracle, a continuously updated QA benchmark leveraging daily news to evaluate the temporal generalization and future prediction capabilities of LLMs. Our experiments reveal that while LLMs maintain a degree of predictive power over future events, their prediction accuracy exhibits a significant smooth decline over time. Although RAG mitigates the effect of outdated knowledge, a strong and noticeable decline remains. Our findings in the gold article setting further emphasize the importance of disentangling missing knowledge from the lack of up-to-date representations. In the future, alongside maintaining Daily Oracle, we plan to incorporate a broader range of models and explore how continuous pre-training and efficient adaptation can address the performance degradation challenges presented in our work.
Impact Statement
Daily Oracle serves as an up-to-date, continuous evaluation benchmark for assessing AI models’ forecasting accuracy and temporal generalization. These abilities have broad applications in areas such as finance, healthcare, and policy-making. The performance drop we observe highlights the risk of outdated knowledge and the importance of continuous model updates to keep AI systems reliable. In the future, by assessing continuous model update strategies in Daily Oracle, the broader ML community can gain valuable insights into how to maintain AI systems that are relevant and well-informed on recent and upcoming events.
@@ Acknowledgment
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. (No. RS-2024-00469482 & RS-2024-00509279) We also thank the Microsoft Accelerating Foundation Models Research program for providing Azure cloud compute credits for the LLM APIs. The compute was also supported by the NYU High Performance Computing resources, services, and staff expertise.
@@ Appendix A Dataset Details
A.1 Illustration of the dataset construction process.
Figure 7 shows how Daily Oracle is automatically generated as discussed in Section 3.1.
Figure 7: Data Construction Process of Daily Oracle.
A.2 Details for Article Selection
@@
We select daily articles that generate the QA pairs in two ways: (1) Random Selection: We randomly sample three articles each day. (2) Hot Topic Selection: To better capture daily events and reduce noise, we select three articles from the top three hot topics of the day. We identify these hot topics by applying the density-based clustering algorithm DBSCAN (Ester et al., 1996) to the new articles based on TF-IDF (Term Frequency-Inverse Document Frequency) representations, forming clusters of news articles for each day. We filter out chaotic clusters by removing those with low average in-cluster cosine similarity scores, which typically correspond to clusters containing a large number of diverse articles. The top three clusters, determined by size, are assumed to represent the most discussed events, i.e. hot topics, since larger clusters indicate more articles covering the same event. One article is picked randomly from each of the top three clusters.
A.3 QA Filtering Principles
During the design stage of QA pair generation, we manually review the questions and identify seven key criteria to ensure the QA pairs qualify as valid forecasting questions. These principles guide the QA filtering step in the data construction process:
  1. (1)
    Correctness of Answers: The answer must be factually accurate and fully aligned with the information in the given article.
  2. (2)
    Non-answerability Before the Publication Date: Since we treat the article’s publication date as the question’s resolution date, the question should not be definitively answerable based on information available before the article’s publication.
  3. (3)
    Absence of Information Leakage: Questions must avoid revealing information that became known only after the article’s publication, maintaining fairness for pre-publication evaluation.
  4. (4)
    Objectivity: Both questions and answers must rely on objective facts, avoiding subjective ideas from the authors.
  5. (5)
    Inclusion of a Clear Temporal Element: Questions must contain a specific and clear reference to time, avoiding vague phrases like “in the future” or “soon.”
  6. (6)
    Public Interest: The questions should address topics of broad public concern.
  7. (7)
    Non-obviousness of the Answer: The answer should not be immediately predictable from the question and must provide new or non-trivial insights.
A.4 Details for Human Evaluation
We assess the quality of our dataset by evaluating the effectiveness of our LLM-based evaluator in the QA Filtering step. Four human annotators independently review a randomly sampled subset of Daily Oracle, consisting of 30 TF and 30 MC QA pairs. They follow the same instructions used to prompt the LLM and evaluate each QA pair based on the seven filtering principles listed in Appendix A.3.
Table 4 presents the inter-rater agreement among human annotators and the agreement between human and LLM evaluators. The average Fleiss’ Kappa of 0.26 indicates fair agreement among annotators. Among the seven principles, Objectivity exhibits the highest agreement (0.66), while Non-Answerability Before the Publication Date has the lowest (0.02).
Comparing human-assigned and LLM-assigned scores, the exact-match accuracy between the human consensus and LLM evaluations averages 89.52% across the seven principles, showing the effectiveness of our LLM-based filtering method. Non-Answerability Before the Publication Date shows the lowest agreement (83.33% accuracy), suggesting it is the most challenging principle for both humans and the LLM to evaluate consistently.
Table 4: The inter-rater agreement among four human annotators is evaluated using Fleiss’ Kappa, while the agreement between human and LLM evaluators is measured through accuracy scores. We report metrics across seven QA Filtering principles using a sample of 60 randomly selected QA pairs.
Human Agreement Human vs. LLM
Metric Fleiss’ Kappa Accuracy (%)
Correctness of Answers 0.11 96.67
Non-answerability Before the Publication Date 0.02 83.33
Absence of Information Leakage 0.33 86.67
Objectivity 0.66 98.33
Inclusion of a Clear Temporal Element 0.21 90.00
Public Interest 0.18 88.33
Non-obviousness of the Answer 0.29 83.33
Average 0.26 89.52
A.5 Distribution of Question Categories Over Time
In Figure 8, we provide the distributions of question categories for both TF and MC questions.
Figure 8: Question category distribution over time.
@@ B.1 Baseline Models Information
B.2 Refusal Rates
Although the models are prompted to provide definitive answers rather than responding like “I cannot predict the future,” some models still occasionally refuse to do so. Figures 9(b), 10(b), 11(b) show the refusal rates for the closed-book, constraint open-book, and gold article settings. In closed-book evaluation (Figure 9(b)), we can see that the refusal rates increase throughout the time for Mistral-7B in TF questions and Mixtral-8x7B in both TF and MC questions. Additionally, these two models exhibit notable refusal rates, with approximately 10–30% on TF questions and 1.5–8% on MC questions, resulting in their closed-book performances dropping below the random baseline of 50% in certain months, as shown in Figure 3. In comparison, Qwen-2-7B and Gemma-2-2B show relatively low refusal rates—$<$5% for TF and $<$2% for MC—while all other models have near-zero refusal rates for TF and $<$1% for MC.
The refusal behavior is likely influenced by alignment techniques, which discourage uncertain responses in the post-training stage. Although refusal rates contribute to lower accuracies for certain models, our results show that performance degradation trends persist even when refusals are excluded (Figures 9(a), 10(a), 11(a)). We consider refusal to answer an indicator of performance limitations in forecasting tasks, as it reflects the model’s lack of actionable knowledge. When models are supplied with more up-to-date and relevant information, their refusal rates decrease (Figure 10(b)). This suggests that refusal is one example of the broader challenge of temporal generalization and reinforces the need for continual model updates or improved external knowledge integration.
Figure 9: Accuracy excluding refusal rates and refusal rates under the closed-book setting. We plot the 5-month moving average refusal rates for TF and MC questions across different LLMs. We count refusal cases as incorrect both to maintain comparability across models and because failing to provide an answer when a prediction is expected represents an unsatisfactory outcome from the user’s perspective.
Figure 10: Accuracy excluding refusal rates and refusal rates for Mixtral-8x7B under the constrained open‑book setting. (b) shows that the open-book refusal rate (blue curves) is lower than in the closed-book setting (cyan curve), indicating that access to more up-to-date and relevant information reduces model refusals.
Figure 11: Accuracy excluding refusal rates and refusal rates under the gold article setting.
@@ B.2 Refusal Rates
To more effectively illustrate the trends of other models at a suitable scale, we display GPT-3.5’s performance in the gold article setting separately. As shown in Figure 12, this outdated model performs relatively poorly throughout. While its accuracy could improve with chain-of-thought prompting (Wei et al., 2022), we report its performance using the same prompt format as the other models for consistency in comparison. Nevertheless, the degrading trend can still be observed.
@@ B.4 More Results in the Constraint Open-Book Setting
Figures 13, 14, 15, 16, 17, 18, and 19 show the constrained open-book evaluation results for more models. Similar patterns are observed as discussed in Section 4.2. Specifically, for Claude-3.5-Sonnet, the constrained open-book performance lags behind its closed-book performance, likely because it already has robust representations of world events, suggesting that irrelevant or confounding retrieved information may degrade performance. This highlights the need for more careful RAG design in models that already possess robust world knowledge. GPT-3.5 is not included in the constrained open-book setting due to its unexpectedly poor performance in the gold article setting (Figure 12) and budget limitations. Additionally, due to budget constraints, open-book evaluations of proprietary LLMs (Claude-3.5-Sonnet, GPT-3.5, GPT-4) are conducted only up to September 2024, whereas other LLMs are evaluated through December 2024.
@@
B.5 An Example of Evaluating LLMs Under Different Settings
Figure 20 presents a case study demonstrating how Mixtral-8x7B responds to a question under different experimental settings. The model provides an incorrect answer in the closed-book setting. However, when supplemented with retrieved relevant articles or the gold article, it produces the correct answer.
{mdframed}
[News Article]
Title: DFL leaders, Minneapolis announce deal on rideshare pay, but Lyft and Uber say they will leave
Publishing Date: 2024-05-06
Text: ST. PAUL, Minn. — Key lawmakers and DFL legislative leaders on Monday announced that they have a compromise on statewide rideshare regulations in Minnesota, but Uber and Lyft quickly rejected the plan and said that they will still exit the market if the proposal becomes law.
The amendment, which will be heard in a House committee Tuesday, includes a minimum wage pay rate of \$1.27 per mile and 49 cents per minute. Leaders said they worked with the Minneapolis City Council on finding compromise, but the companies were not included in the negotiations.
The proposal is lower than the Minneapolis rate of \$1.40 per mile and \$0.49 per minute, but higher than study done by the Minnesota Department of Labor and Industry, which determined \$0.89 mile and \$0.49 for the Twin Cities metro would bring driver pay up to the cityś minimum wage of \$15.57 an hour.
“This was a really big move for the Minneapolis City Council. This is coming down 13 cents from the rate that they passed. And this is consistent with the rates that Uber and Lyft are operating with and other states and other jurisdictions,” House Majority Leader Jamie Long, DFL-Minneapolis, who was a key negotiator, told reporters. “We know that the companies have said the same thing in other states. They’ve said that they were going to leave at certain rates, and they’ve stayed and been able to operate profitably.”
Uber and Lyft quickly responded that they will still plan to leave Minneapolis — as they said would happen on July 1, when the cityś ordinance is effective — and the state if this becomes law.
  [Generated QA Pairs]
Question:
What will be the response from Uber and Lyft if the proposed rideshare regulations become law in Minnesota by May 2024?
Answer: (a) They will plan to leave the market.
Misleading Choices:
(b) They will stay and adjust their pricing to meet regulations.
(c) They will collaborate to develop an alternative service model
(d) They will reduce their services rather than exit.
  [Closed-Book Setting]
  • The question is asked directly to Mixtral-8x7B.
  • Model response: (wrong) “(d) They will reduce their services rather than exit. This is because leaving the market entirely can be a drastic move and may not be the first response to new regulations. Reducing services is a more likely response, as it allows rideshare companies to comply with the new laws while still maintaining a presence in the market. Additionally, they may also try to negotiate or influence the regulations before they become law.”
[Constraint-Open Book Setting]
  • The RAG-cutoff refers to the latest accessible date for retrieving articles. With RAG-cutoff = 2024-03-01, we extracted the top 5 relevant articles (here, only the titles are displayed, but the articles are included in the evaluation prompt):
    • (2023-05-25) Uber Threatens To Cut Most Service In Minnesota If Minimum Pay Law Takes Effect
    • (2023-05-25) Minnesota Governor Vetoes Rideshare Minimum Pay Bill After Uber Threatens To Leave State
    • (2023-08-09) Minneapolis City Council moving towards minimum wage for rideshare drivers
    • (2024-02-27) Minneapolis City Council to consider rideshare minimum wage ordinance Tuesday
    • (2024-01-11) Some Lyft, Uber drivers to strike Thursday at MSP Airport
  • Model response: (correct) “(a) They will plan to leave the market.”
[Gold Article Setting]
  • In the gold article setting, models are provided direct access to the gold article, from which the question is generated. The question then is treated as a reading comprehensive question rather than a forecasting question.
  • Model response: (correct) “(a) They will plan to leave the market.”
Figure 20: An example of evaluating Mixtral-8x7B across different settings.
Appendix C Comparing LLM-Generated and Forecasting Market Datasets
Online forecasting markets such as Metaculus and Polymarket allow users to submit questions and predict the outcomes of future events. A natural question arises: why do we focus on LLM-generated questions rather than sourcing from these markets, as the choice in prior work (Zou et al., 2022; Halawi et al., 2024; Karger et al., 2025)?
To answer this, we analyze the dataset from Halawi et al. (2024), which compiled 50,343 raw questions from five forecasting platforms, of which 21,149 were resolved. Of these, 82.64% are TF questions, 13.36% are MC questions, and the rest are free-response or numerical. After their quality filtering, only 5,516 TF questions remained. We observe that, due to a high proportion of low-quality questions in the raw data, sparse coverage in earlier years, and inconsistent distribution over time, performance trends derived from this dataset are substantially more volatile and harder to interpret than those based on our LLM-generated dataset.
Lower Quality in Raw Questions.
Manual inspection confirms that a substantial portion of the raw dataset consists of low-quality questions, as also noted by the original authors. Examples include: “Will I have a chess.com rating of $>$1300 …?” (personal), “Will Jamaica beat Mexico?” (missing a time frame), and “Are there more disadvantages in AI than advantages?” (ill-defined). From a random sample of 50 questions, only 28% were well-defined. Specifically, 26% lacked a clear time element, 20% were overly personal, and 26% were ill-defined. Importantly, only 5,516 out of 17,477 resolved TF questions were retained after their filtering—an acceptance rate of just 32%, which aligns with our own quality assessments.
Limited Early-Year Coverage.
Figure 21 (left) shows that the coverage before October 2022 is sparse, averaging only  40 raw and  26 filtered questions per month. This scarcity limits the feasibility of longitudinal trend analysis, especially across earlier model pre-training cutoffs. In contrast, our method supports high scalability and retrospective generation, allowing for uniform coverage across the full time range.
Figure 21: Dataset size and closed-book evaluation accuracy over time on the forecasting market dataset.
Harder-to-Discern Trends.
To evaluate the impact of using forecasting market questions in our study, we run a closed-book evaluation on TF questions from both the raw dataset (16,089 questions) and the filtered subset (4,572 questions), starting from 2020-01 (the same start date as our dataset). Notably, the original data is imbalanced, with 61.03% “No” answers in the raw set and 64.28% in the filtered set. After balancing, we retain 12,438 questions from the raw data and 3,232 from the filtered set. As shown in Figure 21 (right), neither the raw nor filtered datasets reveal a clear performance trend—model accuracy fluctuates significantly over time. We believe this is due to several factors:
  • Lower data quality: Approximately 70% of raw questions exhibit quality issues. While the overall dataset sizes are comparable (13,744 in ours vs. 12,438 in the raw market dataset), the difference in quality introduces additional noise, making trends harder to detect.
  • Limited early coverage: Even within the filtered dataset, sparse early coverage and inconsistent monthly volume increase variance and reduce the reliability of time-based trends.
  • Confounding factors: We argue that human-submitted questions introduce more confounding factors than automatically generated ones. Figure 22 shows the distribution of data sources and question categories varies significantly across time (e.g. more sports-related questions in later periods). Human-written questions also may differ widely in style and difficulty, making them harder to control for consistency. In contrast, as shown in Figure 8, our dataset maintains relatively stable distributions over time.
Figure 22: Distribution shift of the filtered forecasting market dataset.
Thus, while we do not claim that LLM-generated questions are of inherently higher quality, we argue that our dataset is better suited for analyzing performance trends over time, due to its scalability, stylistic uniformity, stable category distribution, and reduced susceptibility to human-authored confounders. Moreover, if one sources questions from forecasting markets, the dataset update frequency is dependent on whether people are still actively submitting high-quality forecasting questions to the platform. In contrast, our approach enables daily updates and more comprehensive event coverage, making it a valuable complement to human-curated forecasting benchmarks.
Appendix D Prompts
@@
All the prompts we use are shown in this section. The QA generation prompts and evaluation prompts are adapted from Zhang et al. (2024), and the prompt to categorize our generated questions is taken from Halawi et al. (2024).
@@ 2. Please provide one keypoint that summarizes the new event from the article with the following rules:
Figure 23: Prompt in the Article Summary step, adapted from Zhang et al. (2024).
@@
Figure 24: Prompt in the QA Generation step (part 1), adapted from Zhang et al. (2024).
@@
Figure 25: Prompt in the QA Generation step (part 2), adapted from Zhang et al. (2024).
@@
Figure 26: Prompt in the Misleading Choices Generation step, adapted from Zhang et al. (2024).
@@
Answer: New York officials will implement increased safety precautions for the event.
Reasoning: The answer is straightforward and expected, as it directly restates the premise of the question without providing any new or specific details on how the safety precautions will be implemented or what they might entail.
@@
Figure 31: Closed-book evaluation prompt for TF questions, adapted from Zhang et al. (2024).
@@
Figure 32: Closed-book evaluation prompt for MC questions, adapted from Zhang et al. (2024).
@@
Figure 33: Constrained open-book evaluation prompt for TF questions, adapted from Zhang et al. (2024).
@@
Figure 34: Constrained open-book evaluation prompt for MC questions, adapted from Zhang et al. (2024).
@@
Figure 35: Gold article evaluation prompt for TF questions, adapted from Zhang et al. (2024).
@@
Figure 36: Gold article evaluation prompt for MC questions, adapted from Zhang et al. (2024).
@@
Figure 37: Prompt to categorize the generated questions, taken from Halawi et al. (2024).